# Balancing Act: Distribution-Guided Debiasing in Diffusion Models

### v1

@@ 4.2 Implementation Details
Training h-space classifiers. We start with creating a paired dataset $\mathcal{D}^{h}_{clf}$ of h-vector and attribute labels. Specifically, we take a subset $\mathcal{D}$ of the CelebA-HQ [15] dataset and obtain attribute labels for the images using the pretrained attribute classifier $\mathcal{C}_{a}$. Next, we embed image $\mathbf{I^{i}}\in\mathbf{D}$ to obtain the corresponding h-space representation $\mathbf{\mathcal{H}^{i}}=\{\mathbf{h^{i}_{t}}\}^{t=T}_{t=0}$ using DDIM [42] inversion. This yields a labelled dataset $\mathcal{D}^{h}_{clf}$ with pairs $(\mathcal{H}^{i},\mathbf{y^{i}})$, where $\mathbf{y^{i}}=\mathcal{C}_{a}(I^{i})$ is the predicted attribute label for image $\mathbf{I^{i}}$ (e.g. male / female). Next, we train h-space attribute classifiers $\mathcal{C}^{h}_{a}(\mathbf{h_{t}},t)$ as a linear head over $\mathbf{h_{t}}$ and conditioned on time $t$ (ignored sample index $i$ for brevity). These obtained classifiers generate high-accuracy attribute predictions, as shown in Fig. 3. Further details about the dataset and classifiers are provided in a supplementary document.
Figure 5: Balancing generated data across gender and eyeglass attributes with various guidance strategies. Minority group $male$ and $eyeglasses$ are marked. Image-space Universal Guidance and Sample Guidance results generate imbalanced and poor-quality images. For gender, Sample Guidance is able to generate only a few males while maintaining quality. The proposed Distribution Guidance balances (close to the ratio of $0.50:0.50$) both the attributes and generates high-quality images.
@@
Diffusion Model Architecture. We evaluated our proposed debiasing method for fair generation from two state-of-the-art pretrained DMs: a) an unconditional DM, P2  [4], trained on CelebA-HQ dataset, and b) a text conditional DM StableDiffusion v1.5  [33] trained on the LAION [40] dataset. Both these models have exceptional image generation quality; however, they have significant bias concerning the sensitive face attributes, as shown in the following subsections.
4.3 Evaluation and Comparisons.
We compare our proposed method against two techniques for guidance-based generations for DMs and a state-of-the-art sampling-based technique for debiasing generative models, MagNet [12].
@@
Universal guidance [1] performs guidance in the image space but uses pretrained classifiers trained on the clean image. This resolves the additional requirement of training image classifiers on noisy images. The key idea is to use the DDIM scheduler and predict the approximation of $\mathbf{\hat{x_{0}}}$ from noise image $\mathbf{x_{t}}$, and pass it through pretrained image classifier $\mathcal{C}_{A}$. However, this process has two shortcomings. Firstly, it is slow as we have to backpropagate the gradients from the image classifier. Experimentally, we found the h-space guidance to be $7$ times faster than the image guidance. Secondly, it performs poorly at the early stage due to an inaccurate approximation of $\mathbf{x_{0}}$. We used the same training set of $|\mathcal{D}_{clf}|=2K$ to train the h-space and image space attribute classifiers for a fair comparison.
@@
Gender Race Eyeglasses Metric FD FID FD FID FD FID Random Sampling 0.178 54.59 0.334 60.01 0.251 75.21 Universal Guidance  [1] 0.193 52.10 0.377 93.42 0.189 64.55 H-Sample Guidance 0.113 51.46 0.184 56.53 0.118 57.63 H-Distribution Guidance 0.049 50.27 0.113 52.38 0.014 51.78 Random sampling* 0.307 112.28 0.463 123.97 0.276 117.83 Magnet* 0.267 91.15 0.454 97.05 0.281 106.55
Quantitative evaluation. We evaluate our debiasing method for the single attribute case by generating balanced distributions of individual sensitive attributes - gender, eyeglasses, and race, in Tab. 1. For the scope of this paper, we consider all these attributes as binary, and hence expect a $1:1$ ratio of the sensitive attributes in the synthesized images (for e.g. $0.50$ fraction of males and $0.50$ fraction of females in case of gender). Specifically, we generate $10K$ images from each method per attribute and compute the metrics defined in Sec. 4.1. The proposed guidance method outperforms all the methods in terms of visual quality as measured by FID and bias generation measured by FD. Sample guidance achieves comparable FID for gender; however, higher FD indicates inferior debiasing. The tradeoff between FD and FID is discussed in detail in subsection 4.5. This supports our thesis that the distribution guidance provides enough flexibility during generation, resulting in high-quality outputs even with a high guidance scale.
Multiple attributes. Additionally, we apply our method for debiasing multiple attributes simultaneously in Tab. 2. Specifically, given two reference distributions $\mathbf{p^{a_{1}}_{ref}}$ and $\mathbf{p^{a_{2}}_{ref}}$, we add guidance from two pretrained attribute distribution predictors $g_{\psi_{1}}$ and $g_{\psi_{2}}$. For this experiment, we define both the reference distribution as uniform for each attribute ($50\%$-$50\%$ splits). The generated results follow the reference distribution, resulting in a balanced generation across attributes.
@@
Gender + Race Eyeglasses + Race Gender + Eyeglasses Metric FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ Random Sampling 0.256 60.68 0.292 89.14 0.214 70.97 Universal Guidance 0.283 71.84 0.264 91.54 0.157 80.57 H-Sample Guidance 0.241 59.78 0.135 67.87 0.079 52.03 H-Distribution Guidance 0.075 49.91 0.101 57.46 0.057 47.45
Comparison with MagNet [12]: MagNet is an unsupervised method enabling fair sampling from a pretrained model. They propose a method for uniform sampling on the image manifold to generate under-represented groups equally. We generated results by MagNet sampling from a StyleGAN2 [17] model trained on the FFHQ [16] dataset from their official codebase. We compare our method with MagNet [12] for the task of fair generation in Tab. 1. We used MagNet to generate $10K$ samples and compute the FD score. As the base model for MagNet is StyleGAN2, we cannot compare FID directly with our DM debiased results. However, we can observe that MagNet achieves similar FD scores to its baseline model and is ineffective in debiasing the three sensitive attributes: gender, race, and eyeglasses. However, our distribution guidance significantly improves the FD metric from the baseline DM model.
Qualitative evaluation. We present qualitative results for attribute balancing with respect to gender and eyeglasses in Fig. 5. We randomly sample $20$ starting noise and use individual guidance methods for debiasing. Without guidance, the DM mostly generates female faces for the gender attribute. Sample guidance and universal guidance, although increasing the number of males, collapses some images. On the other hand, Distribution guidance generates an almost equal number of males and females without affecting the generation quality. Moreover, the DM hardly generates images with eyeglasses for sample and universal guidance and also without any guidance. Our method, on the contrary, leads to highly balanced generations.
@@ 4.4 Generating Imbalanced Distributions
We tested our distribution guidance for generating imbalanced attribute distribution by providing skewed $\mathbf{p^{a}_{ref}}$ i) $0.20$ female and $0.80$ male ii) $0.10$ white race and $0.90$ black race. These two settings are extremely challenging given that male and black race are minority groups in the dataset. We present qualitative results for both the settings in Fig.6, where our distribution guidance is able to generate the defined distribution with majority males. Note, we have binarized the race attribute as black and whites for simplicity, and hence in the generation brown race is considered under black category. We report quantitative metrics in Tab. 3, where the proposed method can achieve good FD scores with the same FID.
@@
Table 3: Distribution guidance for imbalanced generation
0.20F - 0.80M 0.10W - 0.90B Method FD FID FD FID Random Sampling 0.478 72.26 0.734 77.63 H-Distribution Sampling 0.168 51.65 0.325 53.80
@@ 4.5 Ablations
Guidance Strength $\gamma$. We ablate over guidance strength parameter $\gamma$ for both the distribution guidance and sample guidance in Fig. 7. Increasing $\gamma$ reduces the bias (better FD) at the cost of inferior image quality (increased FID). Guidance strength $1500$ achieved a good tradeoff between image quality and balancing attributes. Further, as compared to sample guidance, the distribution guidance achieves a better tradeoff for all the guidance strengths.
Figure 7: Ablation over guidance strength $\gamma$
Batch size. As we approximate $\mathbf{p^{a}_{\theta}}$ with a an estimate $\mathbf{\hat{p}^{a}_{\theta}}$ over a batch of size $N$, we ablate over different values of $N$ in Tab. 4. Intuitively, using a larger batch size yields a better estimate of $\mathbf{p^{a}_{\theta}}$. We have found that the $N=100$ works best for all our experiments.
@@
Batch size 10 25 50 75 100 FD $\downarrow$ 0.062 0.059 0.049 0.059 0.046 FID $\downarrow$ 50.54 51.51 50.86 50.98 51.64
Figure 8: Debiasing results on stable diffusion for gender attribute. The proposed distribution guidance can balance the gender attributes in the neutral prompt and generalizes well to prompts involving other professions e.g. firefighter, doctor.
4.6 Experiments on Stable Diffusion
We implemented our distribution guidance technique for debiasing a text-to-image generation model, Stable Diffusion (SD) v1.5 [33] concerning the gender attribute in this subsection. First, we generate a dataset $\mathcal{D}$, of $10K$ images from SD, with prompts ‘a photo of a male’ and ‘a photo of a female’ to generate a labeled dataset for training h-space classifiers. If we use the same CelebA-HQ dataset to train the h-space classifier, the guidance is ineffective due to a significant domain shift from the SD generations. Next, we obtain the corresponding labeled dataset $\mathcal{D}^{h}_{clf}$ in the h-space and train a gender classifier in the h-space. We used the trained classifier for distribution guidance following Sec. 3.4 to generate images with balanced gender.
It is observed that SD increases the gender bias when queried to generate certain professions (e.g. male and doctor) [52]. To this end, we implement our distribution guidance along the prompts ‘a photo of a doctor’ and ‘a photo of a firefighter’ to evaluate the effectiveness in this challenging setting. The qualitative results are shown in Fig.8, and the quantitative results are reported in Tab. 5. We can observe, as compared to the baseline (random generation without guidance), our method generates fairer gender distribution with both neutral prompts and prompts mentioning professions (e.g. ‘a photo of a firefighter’). Additional results on text-to-image models are present in supp. material.
@@ 4.6 Experiments on Stable Diffusion
Method Gender Doctor Firefighter FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ Random Sampling 0.317 72.37 0.355 70.11 0.235 71.86 H-Sample Guidance 0.026 70.96 0.021 68.43 0.097 70.42 H-Distribution Guidance 0.024 70.69 0.015 67.36 0.093 69.41
@@ 4.7 Class-imbalance in attribute classification
We explore an important downstream application of our proposed approach in balancing minority classes by augmenting the under-represented classes with generated data. Specifically, we train a race classifier (labels obtained from $\mathcal{C}_{\mathcal{A}}$) on the CelebA-HQ [14] dataset. The race classifier is an ImageNet [7]-pretrained ResNet-18 [9] encoder, followed by 2 MLP layers and a classifier. We manually oversample the Whites and undersample the Blacks in the training dataset such that the imbalanced dataset consists of $10k$ samples of white people and $1k$ black people (we keep the genders balanced within a race class). Consequently, the model performs poorly on the minority class (i.e., Black) due to under-representation. Next, we augment (class-balanced) the training data by generating samples whose distribution is inversely proportional to the class counts in the training set to increase images of minority classes using our distribution guidance approach. This adds 9k images of only Black race, and the classifier trained on this balanced data performs significantly better (Tab. 6). Even when gender is balanced in both the classes, we observe a significant disparity in the accuracies for Male and Female Black samples in the vanilla classifier. However, our proposed method helps reduce the accuracy gap between Black males and Black females. This shows a potential application in generating ’class’ balanced datasets to train models for other downstream tasks, which can also mitigate bias.

### v2

@@ 4.2 Implementation Details
Training h-space classifiers. We start with creating a paired dataset $\mathcal{D}^{h}_{clf}$ of h-vector and attribute labels. Specifically, we take a subset $\mathcal{D}$ of the CelebA-HQ [20] dataset and obtain attribute labels for the images using the pretrained attribute classifier $\mathcal{C}_{a}$. Next, we embed image $\mathbf{I^{i}}\in\mathbf{D}$ to obtain the corresponding h-space representation $\mathbf{\mathcal{H}^{i}}=\{\mathbf{h^{i}_{t}}\}^{t=T}_{t=0}$ using DDIM [51] inversion. This yields a labelled dataset $\mathcal{D}^{h}_{clf}$ with pairs $(\mathcal{H}^{i},\mathbf{y^{i}})$, where $\mathbf{y^{i}}=\mathcal{C}_{a}(I^{i})$ is the predicted attribute label for image $\mathbf{I^{i}}$ (e.g. male / female). Next, we train h-space attribute classifiers $\mathcal{C}^{h}_{a}(\mathbf{h_{t}},t)$ as a linear head over $\mathbf{h_{t}}$ and conditioned on time $t$ (ignored sample index $i$ for brevity). These obtained classifiers generate high-accuracy attribute predictions, as shown in Fig. 5. Further details about the dataset and classifiers are provided in Sec.D of supplement.
Figure 6: Balancing generated data across gender and eyeglass attributes with various guidance strategies. Minority group $male$ and $eyeglasses$ are marked. Image-space Universal Guidance and Sample Guidance results generate imbalanced and poor-quality images. For gender, Sample Guidance is able to generate only a few males while maintaining quality. Latent Editing produces collapsed images for gender and fails to generate eyeglasses. The proposed Distribution Guidance balances (close to the ratio of $0.50:0.50$) both the attributes and generates high-quality images.
@@
Diffusion Model Architecture. We evaluated our proposed debiasing method on two state-of-the-art pretrained DMs: a) an unconditional DM, P2 [5], trained on CelebA-HQ dataset, and b) a text conditional DM StableDiffusion v1.5 [40] trained on the LAION [48] dataset. Both these models have exceptional image generation quality; however, they have significant bias concerning the sensitive face attributes, as shown in the following subsections.
4.3 Baselines
We compare our proposed method against two techniques for guidance-based generations for DMs, one Latent-based editing method [23], and a state-of-the-art sampling-based technique for debiasing generative models, MagNet [14].
@@ 4.3 Baselines
Universal guidance [1] performs guidance in the image space but uses pretrained classifiers trained on the clean image. This resolves the additional requirement of training image classifiers on noisy images. The key idea is to use the DDIM scheduler and predict the approximation of $\mathbf{\hat{x_{0}}}$ from noise image $\mathbf{x_{t}}$, and pass it through pretrained image classifier $\mathcal{C}_{A}$. However, this process has two shortcomings: it is slow as it backpropagates the gradients from the image classifier (Experimentally, we found it to be $7$ times slower than h-space guidance) and it performs poorly at the early stage due to an inaccurate approximation of $\mathbf{x_{0}}$. We use two settings where we vary the number of images in the training set of the image space attribute classifier: (1) Using the same training set $|\mathcal{D}_{clf}|=2K$ as the h-space classifier and (2) Using the entire CelebA-HQ dataset, i.e. 30k images.
Latent-based Editing [23] generates images with a specific set of attributes. Such a technique is popularly used in debiasing GANs [37, 17] because of the well-known disentangled latent spaces of GANs that allow for such edits. Recent works have shown  [23] that similar semantic control is also present in the h-space of DMs and can be used for latent-based editing. We capitalize on this finding and perform latent editing in the h-space to generate images of desired attributes for fair generation.
MagNet [14] is an unsupervised method enabling fair sampling from a pretrained model. They propose a method for uniform sampling on the image manifold to generate under-represented groups equally. We generated results by MagNet sampling from a StyleGAN2 [22] model trained on the FFHQ [21] dataset from their official codebase. Notably, as the base model for MagNet is StyleGAN2, we cannot directly compare FIDs with our DM debiased results and report random generations from StyleGAN2 as a reference.
4.4 Main Results
We first present the quantitative and qualitative results on debiasing single binary attributes. Second, we debias multiple attributes simultaneously. We finally present the case of debiasing attributes with multi-class labels.
@@ 4.4 Main Results
Gender Race Eyeglasses Method FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ Random Sampling 0.178 54.59 0.334 60.01 0.251 75.21 Universal Guidance (2k) [1] 0.193 52.10 0.377 93.42 0.189 64.55 Universal Guidance (30k) [1] 0.127 48.94 0.326 58.52 0.051 78.57 Latent Editing [23] 0.001 37.40 0.214 42.69 0.330 75.04 H-Sample Guidance (ours) 0.113 51.46 0.184 56.53 0.118 57.63 H-Distribution Guidance (ours) 0.049 50.27 0.113 52.38 0.014 51.78 StyleGAN2 - Random sampling 0.307 112.28 0.463 123.97 0.276 117.83 StyleGAN2 - Magnet [14] 0.267 91.15 0.454 97.05 0.281 106.55
Quantitative evaluation. We evaluate our debiasing method for the single attribute case by generating balanced generations of individual sensitive attributes - gender, eyeglasses, and race, in Tab. 1. As these attributes are binary, the synthesized images are expected to have a $1:1$ ratio of the sensitive attributes (for e.g. $0.50$ fraction of males and $0.50$ fraction of females in case of gender). Specifically, we generate $10K$ images from each method per attribute and compute the metrics defined in Sec. 4.1. For most attributes, the proposed guidance method outperforms all the baselines in terms of visual quality as measured by FID and bias metric measured by FD. Although Latent editing has a better FID and FD for gender, on qualitative evaluation (as elaborated in the next section), artifacts are seen in the images (Fig. 6). Moreover, this methods fails to mitigate bias in case of multiple attributes (Tab. 3). Sample guidance achieves comparable FID for gender; however, higher FD indicates inferior debiasing. The tradeoff between FD and FID is discussed in Sec.A.1 of supplement. This supports our thesis that the distribution guidance provides enough flexibility during generation, resulting in high-quality outputs even with a high guidance scale.
Qualitative evaluation. We present results for balancing gender and eyeglasses attributes in Fig. 6. We randomly sample $20$ starting noise and use individual guidance methods for debiasing. Without guidance, the DM mostly generates female faces for the gender attribute. Although latent editing achieved better quantitative metrics, it produces images with artifacts and leads to collapsed results. Although Sample and Universal guidance increase the number of males, some images collapse. On the other hand, distribution guidance generates an almost equal number of males and females without affecting the generation quality. Moreover, all the baselines are not able to generate eyeglasses, whereas our method leads to highly balanced generations.
Multiple attributes. We apply our method for debiasing multiple attributes simultaneously in Tab. 3. Specifically, given two reference distributions $\mathbf{p^{a_{1}}_{ref}}$ and $\mathbf{p^{a_{2}}_{ref}}$, we add guidance from two pretrained attribute distribution predictors $g_{\psi_{1}}$ and $g_{\psi_{2}}$. For this experiment, we define both the reference distribution as uniform for each attribute ($50\%$-$50\%$ splits). The generated results follow the reference, resulting in a balanced generation across attributes. Further analysis is provided in Sec.B of supplement.
@@
Gender + Race Eyeglasses + Race Gender + Eyeglasses Method FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ Random Sampling 0.256 60.68 0.292 89.14 0.214 70.97 Latent Editing [23] 0.124 64.84 0.219 90.63 0.230 74.93 Universal Guidance (2k) [1] 0.283 71.84 0.264 91.54 0.157 80.57 H-Sample Guidance (ours) 0.241 59.78 0.135 67.87 0.079 52.03 H-Distribution Guidance (ours) 0.075 49.91 0.101 57.46 0.057 47.45
Multi Class attributes. We evaluate the efficacy of our approach in balancing multi-class attributes - age and race in Tab. 3. We use FFHQ [21] dataset for both attributes and obtained annotations using pretrained models as ground truth labels are unavailable. For age attribute, we use a pretrained VIT age classifier [39] to produce $3$ classes: Young (${<20}$ yrs), Adult (${20-60}$ yrs) and Old (${>60}$ yrs). For race, we use the Fairface race classifier [16] to obtain $4$ classes: White, Black, Asian, and Indian. Our method successfully debiases multi-class attributes and beats the random and sample guidance in all cases.
@@ 4.5 Generating Imbalanced Distributions
We test our distribution guidance for generating imbalanced attribute distribution by providing skewed $\mathbf{p^{a}_{ref}}$ i) $0.20$ female and $0.80$ male ii) $0.10$ white race and $0.90$ black race. These two settings are extremely challenging given that male and black race are minority groups in the
Table 3: Balanced generation for multi-class attribute
Age (3 classes) Race (4 classes) Method FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ Random Sampling 0.256 60.68 0.292 89.14 H-Sample Guidance 0.124 64.84 0.219 90.63 H-Distribution Guidance 0.283 71.84 0.264 91.54
Table 4: Distribution guidance for imbalanced generation
0.20F - 0.80M 0.10W - 0.90B Method FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ Random Sampling 0.478 72.26 0.734 77.63 H-Distribution Sampling 0.168 51.65 0.325 53.80
dataset. We present qualitative results for both the settings in Fig.7, where our distribution guidance is able to generate the defined distribution with majority males. Note, we have binarized the race attribute as black and whites for simplicity, and hence in the generation, brown race is considered under black category. We report quantitative metrics in Tab. 4, where the proposed method can achieve good FD scores with the same FID. Additional experimental results are tabulated in Sec.B.1 of supplement.
@@ 4.6 Ablations
We provide ablation over the batch size here and that over the guidance scale, h-space classifier architectures and number of training examples in Sec.A of supplement. Batch size. As we approximate $\mathbf{p^{a}_{\theta}}$ with a an estimate $\mathbf{\hat{p}^{a}_{\theta}}$ over a batch of size $N$, we ablate over different values of $N$ in Tab. 5. Intuitively, using a larger batch size yields a better estimate of $\mathbf{p^{a}_{\theta}}$. We have found that the $N=100$ works best for our experiments, balancing both FD and FID. Additionally, our model can also handle low data regime effectively.
@@
Batch size 2 4 8 10 25 50 75 100 125 150 200 FD $\downarrow$ 0.108 0.088 0.073 0.062 0.059 0.049 0.059 0.046 0.052 0.053 0.058 FID $\downarrow$ 60.12 51.86 49.80 50.54 51.51 50.86 50.98 51.64 49.91 49.81 49.74
Figure 8: Debiasing results on stable diffusion for gender. Distribution guidance can balance the gender attributes in prompts involving other professions e.g. firefighter and doctor.
4.7 Debiasing Text-to-Image Diffusion Models
We implemented our distribution guidance technique for debiasing a text-to-image generation model, Stable Diffusion (SD) v1.5 [40] concerning the gender attribute in this subsection. We provide results on other attributes and on mitigating spurious correlation on WaterBirds [43] generation in Sec.C of supplement. First, we generate a dataset $\mathcal{D}$, of $10K$ images from SD, with prompts ‘a photo of a male’ and ‘a photo of a female’ to generate a labeled dataset for training h-space classifiers. If the h-space classifier is trained on CelebA-HQ dataset, guidance is ineffective due to a significant domain shift from the SD generations. Next, we obtain the corresponding labeled dataset $\mathcal{D}^{h}_{clf}$ in the h-space and train a gender classifier in the h-spaceh-space. We used the trained classifier for distribution guidance following Sec. 3.4 to generate images with balanced gender.
It is observed that SD increases the gender bias when queried to generate certain professions (e.g. male and doctor) [62]. To this end, we implement our distribution guidance along the prompts ‘a photo of a doctor’ and ‘a photo of a firefighter’ to evaluate the effectiveness in this challenging setting. The qualitative results are shown in Fig.8, and the quantitative results are reported in Tab. 6.
Baselines. We compare our method with the following methods. (1) Random sampling from SD  [40]. (2) ITI-Gen  [62] that learns prompt embeddings for each category of the attribute given image reference sets of each category. It then appends these prompts during generation to produce balanced images. (3) Fair Diffusion  [10] that uses a lookup table to recognize the biased concept from the text input and adds scaled attribute expressions to the prompt. Note, these baselines are explicitly designed to debias text-conditioned diffusion model; however, our method can debias both conditional and unconditional diffusion models.
@@ 4.7 Debiasing Text-to-Image Diffusion Models
Method Gender Doctor Firefighter FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ FD $\downarrow$ FID $\downarrow$ Random Sampling 0.317 72.37 0.355 70.11 0.235 71.86 ITI-Gen  [62] 0.049 64.79 0.072 67.81 0.184 70.12 Fair Diffusion  [10] 0.227 71.22 0.035 74.37 0.036 68.33 H-Sample Guidance (ours) 0.026 70.96 0.021 68.43 0.097 70.42 H-Distribution Guidance (ours) 0.024 70.69 0.015 67.36 0.093 69.41
@@ 4.8 Class-imbalance in attribute classification
We explore an important downstream application of our proposed approach in balancing minority classes by augmenting the under-represented classes with generated data. Specifically, we train a race classifier (labels obtained from $\mathcal{C}_{\mathcal{A}}$) on the CelebA-HQ [19] dataset. The race classifier is an ImageNet [8]-pretrained ResNet-18 [11] encoder, followed by 2 MLP layers and a classifier. We manually oversample the Whites and undersample the Blacks in the training dataset such that the imbalanced dataset consists of $10k$ samples of white people and $1k$ black people (we keep the genders balanced within a race class). Consequently, the model performs poorly on the minority class (i.e., Black) due to under-representation. Next, we augment (class-balanced) the training data by generating samples whose distribution is inversely proportional to the class counts in the training set to increase images of minority classes using our distribution guidance approach. This adds 9k images of only Black race, and the classifier trained on this balanced data performs significantly better (Tab. 7). Even when gender is balanced in both the classes, we observe a significant disparity in the accuracies for Male and Female Black samples in the vanilla classifier. However, our proposed method helps reduce the accuracy gap between Black males and Black females. This shows a potential application in generating ‘class’ balanced datasets to train models for other downstream tasks, which can also mitigate bias.